OpenStudy (caozeyuan):

Calculate the pH of a 0.30 M NaF solution. The Ka value for HF is 7.2*10^-4

OpenStudy (caozeyuan):

Please explain the procedure involved in solving this kind of problem in general, medal's given if answer is clear

OpenStudy (abb0t):

Lol we're not here competing for medals you know...

OpenStudy (frostbite):

Totally right abb0t. Okay, we look at the \(\sf F^{-}\) ion and remember it is a weak base. So we are going to look at the reaction: \[\large \sf F^{-} + H _{2}O \to FH + OH ^{-}\] The base dissociation constant become: \[\large \sf K _{b}= \frac{ \left[ FH \right] \times \left[ OH ^{-} \right] }{ \left[ F ^{-} \right] }\] We want to find the \(\sf H^{+}\) concentration so we need to find \(\sf \left[ OH^{-} \right]\). We do so by assuming that the reaction run \(x\) times and we then get: \[\large \sf K _{b}=\frac{\it x \times x }{ \it 0.30-x }\] Solve for \(x\) as \(x=\sf \left[ FH \right]=\left[ OH ^{-} \right]\) When we got \(\sf \left[ OH ^{-} \right]\) we use the constant for the self-ionization of water which is equal to: \[K _{w}=\sf \left[ OH ^{-} \right] \times \left[ H ^{+} \right]=10^{-14} ~ M ^{2} ~ ~ ~ ~ T=25^{\circ}C\] Then use the definition of pH to find pH.

OpenStudy (frostbite):

Oh yea forgot: Start by calculating \(\sf K_{b}\).

OpenStudy (caozeyuan):

@Frostbite , The value of Ka is given! so I think there is no need to calculate the value of Kb. And our Chem teacher mentioned something called " 5% rule " that can be used to discard the "-x" part in "0.30-x". How does that comes in?

OpenStudy (caozeyuan):

sorry it should be HF solution, not NaF. I misread the question last night because I was too tired

OpenStudy (frostbite):

Well, I just answered what you asked. :) About the "5% rule" I don't know what that is. We do have ways to approximate our self through acid/base calculations, but usually it is just better not to do so, as it is more easy and you don't do it wrong.

OpenStudy (anonymous):

A buffer solution may be acidic or basic. Your question involves the acidic buffer. An acidic buffer solution contains a weak acid and its salt. In your question the weak acid is hydrofluoric acid; HF .... Ka = 7.2x10^-4 and its salt is sodium fluoride (NaF) As these formulas imply hyrdofluoric acid is a monoprotic acid. Concentrations of the substances; [HF] = 0.20 M [NaF] = 0.10 M (a) Dissociations: Weak acid ionizes only a small extent. HF(aq) <-----> H+(aq) + F-(aq) 0.20 - x M ...... x M ........ x M Ionic soluble salt ionizes completely. NaF(aq) -----> Na+(aq) + F-(aq) 0.10 M ........... 0.10 M ..... 0.10 M F- is the common ion, but the concentration of F- ion produced by the dissociation of the salt (NaF) will be very large compared to the concentration of F- ion produced by the dissociation of the acid, and hence x is neglected. Ka = [H+][F-] / [HF] This expression in some cases simplified as (since x is neglected) Ka = [H+][SALT] / [ACID] Substituting the values; 7.2x10^-4 = [H+] (0.10) / (0.20) [H+] = (7.2x10^-4 x 0.20) / 0.10 = 1.44x10^-3 Awnser:1.4 x 10-3 M Did that help?